OpenStudy (anonymous):

The curve C is defined by r(t)=<2t, 3sint, 3cost> Find the length of this curve for -4<=t<=4

OpenStudy (amistre64):

\[\int ds\] sounds familiar

OpenStudy (anonymous):

I started some work on this but the answer I was getting was quite messy so i was doubting if it was correct. \[\int\limits_{-4}^{4} \sqrt{4+9\cos^2-9\sin^2}\] \[= \int\limits_{-4}^{4} \sqrt{4+9(1-2\sin^2t)}\] \[ = \int\limits_{-4}^{4} \sqrt{4+9-18\sin^2t}\] \[= \sqrt{13-18 \sin^2t}\]

OpenStudy (amistre64):

would cos(2t) work any better than (1-2sin^2) ?

OpenStudy (anonymous):

scratch that last one missed the integral. so it would be \[\int\limits_{-4}^{4} \sqrt{13-18\sin^2t}\] \[=\frac{ 2 }{ 3 } (13-18\sin^2t)^\frac{ 3 }{ 2 }(13t-18(\frac{ 1 }{ 2 }t-\frac{ 1 }{ 4 } \sin2t))\] then evaluated that last one from -4 to 4 quite messy

OpenStudy (anonymous):

possibly ill retry

OpenStudy (amistre64):

site went out for me ....

OpenStudy (amistre64):

x'^2 + y'^2 + z'^2 there shouldnt be any negative, or subtractions

OpenStudy (anonymous):

ok so carrying through where that changes \[=\int\limits_{-4}^{4} \sqrt{4+9\cos2t}\] \[=\int\limits_{-4}^{4} ( 4+9\cos2t)^\frac{ 1 }{ 2 }\] \[=\frac{ 2 }{ 3 }(4+9\cos2t)^\frac{ 3 }{ 2 }(4t+\frac{ 9 }{ 2 }\sin2t)\] evaluated form -4 to 4

OpenStudy (anonymous):

does that look like it would be a correct answer or is something still off prior to evaluating?

OpenStudy (amistre64):

\[\sqrt{[(2t)']^2+[(3sin(t))']^2+[(3cos(t))']^2}\] \[\sqrt{[2]^2+[3cos(t)]^2+[-3sin(t)]^2}\] \[\sqrt{4+9cos^2(t)+9sin^2(t)}=\sqrt{4+9(cos^2+sin^2)}=\sqrt{13}\]

OpenStudy (anonymous):

oh i see ok i had just found where that was in the work. thanks that is alot better. :)

OpenStudy (amistre64):

:) good luck